A large disc has mass 2kg,radius 0.2m and initial angular velocity 50rad/s and a small disc of mass 4 kg,radius0.1m and initial angular velocity 200rad/s.find the common final angular velocity after the disc are pushed into contact:is the kinetic energy conserved during this process?


1
Expert's answer
2021-03-23T11:16:15-0400

I will solve the question assuming that both the disk are rotating in same direction.

"m_1=2\;kg \\

r_1=0.2 \; m \\

I_1= \frac{1}{2}mr^2 = (0.2)^2= 0.04 \\

m_2=4 \; kg \\

r_2=0.1 \;m \\

I_2= \frac{1}{2}mr^2 = 2 \times (0.1)^2= 0.02"

Let the final angular velocity be w.

Since the external torque on system is 0. The angular momentum of the system will be conserved.

"I_1 \times w_1+I_2 \times w_2 = (I_1+I_2) \times w \\

0.04 \times 50+0.02 \times 200=(0.04+0.02) \times w \\

6=0.06 \times w \\

w=100 \; rad/sec \\

Energy \; initial = \frac{1}{2}I_1 \times w_1^2+\frac{1}{2}I_2 \times w_2^2= 50+400= 450 \\

Energy \; final =\frac{1}{2}I_1 \times w^2+\frac{1}{2}I_2 \times w^2= 300"

Kinetic Energy lost "= 450-300 = 150 \; J" (energy is lost in the work done by kinetic friction)

No the kinetic energy will not be conserved.